OpenStudy (kendricklamar2014):

Is this true? 3^3 * 3^3 > 3^5

OpenStudy (kendricklamar2014):

My answer: True

geerky42 (geerky42):

Correct

OpenStudy (anonymous):

ye

OpenStudy (kendricklamar2014):

thanks

geerky42 (geerky42):

Because 3^3 * 3^3 = 3^6

geerky42 (geerky42):

no problem

OpenStudy (kendricklamar2014):

thats what I thought ^

OpenStudy (kendricklamar2014):

@geerky42 can u help me with one more?

geerky42 (geerky42):

sure

OpenStudy (cwrw238):

laws of indices a^n * a^m = a^(n+m)

OpenStudy (kendricklamar2014):

Write the value of the expression: 3^4/3^4

OpenStudy (kendricklamar2014):

My answer: 1

geerky42 (geerky42):

right again

OpenStudy (kendricklamar2014):

thanks